OpenStudy (anonymous):

If you change the y-intercept in the equation y = x + 7 to y = x + 11, how will its graph change? A. It will increase at a faster rate than before. B. It will increase at a slower rate than before. C. The starting value will increase. D. The starting value will decrease.

OpenStudy (anonymous):

What does "y-intercept" mean if you're looking at a graph?

OpenStudy (anonymous):

im not sure im not good at graphing

OpenStudy (anonymous):

So if you have a graph, the y-intercept and x-intercepts are where the line hits the y and x axes. 

OpenStudy (anonymous):

What can you tell about the value of x at the y intercept?

OpenStudy (anonymous):

like y goes up and down and x goes side to side

OpenStudy (anonymous):

Those are the "x axis" and the "y axis." At the center, where they meet (the origin) is where x and y both are 0.  Straight up the "y axis" the values of y change, but the value of x stays 0 Straight across the "x axis" the values of x change, but y stays 0. And at the y intercept, x = 0. So if we call that the "starting value:" Looking at the given equation, what can you tell me about each value of y given that x = 0? y = x + 7 with x = 0 = ? and y = x + 11 with x = 0 = ?

OpenStudy (anonymous):

ohh okay can I ask if it is C?

OpenStudy (anonymous):

How would you justify that answer if you had to?

OpenStudy (anonymous):

Just with words is fine ^_^

OpenStudy (anonymous):

because it would increase until they meet

OpenStudy (anonymous):

What's "they?" And I'm not sure what you're saying....

OpenStudy (anonymous):

(you are correct, by the way, just wanting to know how you got there)

OpenStudy (anonymous):

(rather, your multiple choice is correct. Your justification isn't making good sense to me.)